OpenStudy (anonymous):

can anyone help me with this 2sintcost - cost = 0 im probably overthinking this

OpenStudy (amistre64):

2 sin(t) cos(t) = sin(2t) when does sin(2t) = cos(t)?

OpenStudy (dumbcow):

factor out the cos

OpenStudy (amistre64):

or...do it the smart way and factor out the ...... lol

OpenStudy (dumbcow):

your way works too

OpenStudy (anonymous):

sint(cost+sint)?

OpenStudy (amistre64):

cos(t)(2sin(t) - 1) = 0 when does cos(t) = 0? 90 and 270 right?

OpenStudy (amistre64):

2sin(t) - 1 = 0 2sin(t) = 1 sin(t) = 1/2...thats a 30 right? and 150?

OpenStudy (anonymous):

k, i knew i had to factor out, just didnt know where to start